my license tag is a 3 digit number.the product of the digit is 216. their sum is 19 and the digits of here and ascending order find the license plate number
Answer by Edwin McCravy(20054)

my license tag is a 3 digit number.the product of the digit is 216. their sum is 19 and the digits of here and ascending order find the license plate number

If the smallest digit were 0 for the product of the digits would be 0.
So 0 is eliminated.

The smallest digit cannot be 1 for the product of the other two digits 
would have to be 216, and that's impossible.  So 1 is eliminated

The smallest digit cannot be 2 as because 216/2 = 108
and so the other two would have to have product 108 and no two digits
have product that large.

If 3 is its smallest digit, then since 216/3 = 72, the other two digits
would have to have product 72, and that is only possible with digits
8 and 9, but 389, is not possible because 3+8+9=20, not 19.

If 4 is its smallest digit, then since 216/4 = 54, the other two digits
would have to have product 54, and that is only possible with digits
6 and 9, 4+6+9=19.   So 469 is an answer since the digits are ascending.

But we need to rule out any other answers.

5 cannot be the smallest digit because 216 is not a multiple of 5.

If 6 is its smallest digit, then since 216/6 = 36, the other two digits
would have to have product 36, and since we've ruled out 4, the only 
possible way is with the other two digits being 6 and 6.  but 666 is not
possible for two reasons, the digits are not ascending and because 6+6+6=18,
not 19.

The only answer is 469.   

Edwin